# Scintillation Properties of PSR B1133+16 Measured with Very Long Baseline Interferometry

### v1

@@ Abstract
The scintillation of pulsars reveals small-scale structure of the interstellar medium. A powerful technique for characterizing the scintillating structures (screens) combines analysis of scintillation arcs and very long baseline interferometry (VLBI). We present the results of a VLBI analysis of the scintillation arcs of PSR B1133+16 from simultaneous observations with Arecibo, VLA, Jodrell Bank, Effelsberg, and Westerbork. Three arcs appear in the data set, all of which appear consistent with being the result of very anisotropic scattering screens. We are able to measure their orientations on the sky, down to uncertainties of $3\arcdeg$ for the two stronger screens, and measure distances, of $140\pm 30$, $180\pm 20$, and $280\pm 50{\rm\,pc}$, consistent with, but substantially more precise than what was inferred previously from annual modulation patterns in the scintillation. Comparing with the differential dust extinction with distance in this direction, the two nearer screens appear associated with the wall of the Local Bubble.
@@ 2.2 Visibility Secondary Cross-Spectra
$\frac{d\Phi_{S_{V}}}{df_{D}}=-\frac{2\pi|\mathbf{b}_{mn}|\cos(\alpha_{mn}-\alpha_{s})}{v_{\textrm{eff},\parallel}}=2\pi\Delta t_{{\rm ISS},mn},$ (8)
where $\alpha_{mn}$ and $\alpha_{s}$ represent the position angles of the baseline $\mathbf{b}_{mn}$ and the scattered images in the $(u,v)$ plane, respectively, and we implicitly defined $\Delta t_{{\rm ISS},mn}$, the time delay between when the scintillation pattern arrives at station $n$ compared to station $m$. One sees that the delay is maximized when the baseline aligns with the line of images, while it becomes zero if it is perpendicular to it: at that orientation, for a strictly linear set of images, the telescopes see identical interference patterns.
@@ 2.3 Intensity Secondary Cross-Spectra
$\frac{d\Phi_{S_{I}}}{df_{D}}=-\frac{2\pi|\mathbf{b}_{mn}|\cos(\alpha_{mn}-\alpha_{s})}{v_{\textrm{eff},\parallel}}=2\pi\Delta t_{ISS,mn}.$ (11)
@@ 2.4 Solving for Screen Parameters
A scintillation screen producing a (nearly) one-dimensional set of scattering images is described by three parameters, $d_{\rm eff}$, $v_{{\rm eff},\parallel}$, and $\alpha_{s}$. These three enter measurables via the curvature $\eta$ and the time delays $\Delta t_{{\rm ISS},mn}$, i.e., via Equations 6, 8, and 11 — all other parameters in those equations are known. Specifically, with at least two time delays (on non-parallel baselines), one can solve for $v_{{\rm eff},\parallel}$ and $\alpha_{s}$, and then combine $v_{{\rm eff},\parallel}$ with the curvature to find $d_{\rm eff}$.
@@ 3 Observations and Data Reduction
We analyze 2 hours of European VLBI Network (EVN) observations of PSR B1133+16 taken on 2019 January 27 (MJD 58510) from 6:29 to 8:29 UTC (PI: Kirsten, project code: GK049E). Data were taken with Arecibo, the Very Large Array (VLA), Jodrell Bank, Effelsberg, Westerbork and Sardinia, but we had to exclude the Sardinia data as these contained no pulsar signal. For phase calibration of the VLA, the observations were interspersed every 26 minutes with a one-minute observation of the source J1143+221. For bandpass calibration, all telescopes observed J1125+2610, using two five-minute scans separated by 80 minutes. The frequency range covered was from 316 MHz to 348 MHz, in two contiguous 16 MHz bands. All telescopes used circular polarizations except for VLA, which was linear. The dual-polarization baseband data were recorded in 2 bit real values, using VDIF format at all telescopes but Arecibo, where Mark 5B was used.
@@ 3.1 Reduction of Visibilities
Calibration of the data was done using CASA (McMullin et al., 2007). For our analysis here, which focuses on the properties of the screens, phase and polarization information is not relevant, so we averaged the eight on-pulse bins together and discarded the cross-polarizations. We flagged bad data and radio frequency interference, combining the flag table output from the EVN pipeline with manual inspection, and then used standard CASA routines for a-priori gain calibration, fringe fitting, bandpass calibration, delay calibration, and amplitude and phase corrections. As a significant portion of the bandpass calibrator observations were flagged or did not include all antennas, we found that the bandpass was flattened more successfully when using the pulsar for calibration directly instead of the calibrator source.
@@ 3.2 Reduction of Intensities
where ${\cal F}$ and ${{\cal F}^{-1}}$ denote the Fourier transform and its inverse, and $\tilde{I}$ and $\tilde{F}$ are the uncorrected Fourier transforms of $I(\nu,t)$ and $F(t)$, respectively. Note that $\tilde{I}_{m}\tilde{I}_{n}^{*}$ is the uncorrected intensity secondary cross-spectrum; since the intensities are real values, it equals $\tilde{I}_{m}(\tau,f_{\rm D})\tilde{I}_{n}(-\tau,-f_{\rm D})$.
@@
Figure 2: Visibility secondary cross-spectra. Top: Phases and intensities for all baselines involving Arecibo (binned by a factor of 4 in $f_{\textrm{D}}$ and 8 in $\tau$). Parabolae for each arc for the best fit arc curvature are overdrawn, with colours following the panels below. Bottom: Phase along the arc, determined by averaging the complex values along delay within a parabolic region around each arc of width $0.58{\rm\,mHz}$. Grey points indicate all measured values, while black ones indicate the ones used for fitting gradients. The best fit is shown with coloured lines, with the uncertainties represented by the surrounding shaded regions.
@@ 4 Scintillation Properties
In Figures 2 and 3, we show the visibility and intensity secondary cross-spectra, with color representing phase. One sees that on the Arecibo-VLA baseline, the inner and middle arc have similar phase behaviour, while on the Arecibo-Jodrell Bank baseline, the phases change in opposite directions along the parabola, showing immediately that these two screens have different orientations. We measure the time delays corresponding to the phase gradients in Section 4.2.
@@ 4.1 Arc Curvatures
To estimate our curvature, we over-plotted parabolae over the secondary spectrum and verified the curvatures by eye. Our uncertainties were determined by eye, by over-plotting a range of curvatures and estimating a valid range. We list the results in Table 1.
@@ 4.2 Phase Gradients
To estimate the time delays along each baseline, we measured the phase gradients as a function of Doppler in the visibility and intensity secondary cross-spectra (see Eqs. 8 and 11). First, we averaged the complex secondary cross-spectra over 4 and 8 bins in Doppler and delay, respectively, i.e., to a resolution of $0.56{\rm\,mHz}$ and $0.25{\rm\,\mu s}$. Then, for each arc, we averaged the complex values along the delay axis within a parabolic mask – centered around the best fit parabola for the given arc and with a small width, $0.58{\rm\,mHz}$ – and determined the phases from those. For the uncertainties of the complex averages, we used the error in the mean, with the noise in the individual points estimated from the spectra far away from the arcs. We show the delay-averaged phases for each arc and baseline together with the binned secondary cross-spectra in Figures 2 and 3.
@@
We tested the influence of our assumptions on our fitted values by repeating the fits using a variety of parabolic masks and fitting regions. Specifically, we varied the arc curvature for the masks within the arc curvature fitting error, the widths from 0.56 to 1.12 mHz, and the minimum and maximum Doppler ranges by ±$\pm 2{\rm\,mHz}$ from those listed above. The fits depended most strongly on the arc curvature of the parabolic mask and had very little dependence on the widths of the parabolic mask or on the Doppler ranges. Overall, though, the fits deviated from each other beyond the nominal fitting error, so we take the standard deviation among the trials as the estimate of our uncertainty. We list the results in Table 1.
@@ 4.3 Phase De-Rotation
As an alternative method to determine the phase gradients in the secondary cross-spectra, we applied phase gradients of opposite sign directly to the secondary cross-spectra, i.e. multiplied by $\exp(-\Phi/f_{\textrm{D}})$. If the applied phase gradient is close to the true phase gradient, the imaginary component of the secondary cross-spectrum vanishes or, equivalently, the real component is maximized. We thus tried to find the best fit phase gradient by first computing the sum of real values within the same masked regions as above for several test phase gradients, and then determining the location of the maximum by fitting a parabola to the results. This method gave consistent results to fitting the slope of the phases with Doppler.
@@ 5 Screen Properties
The screen parameters are found by using Equations 8 and 11 to infer from the time delays the orientation of the line of images and the effective velocity for each screen. For this purpose, for each screen we fit all corresponding time delays simultaneously, derived from both intensity and visibility secondary cross-spectra, for the baselines between Arecibo and the four other antennas. Next, we use Equation 6 to calculate the effective distance from the effective velocity and the measured arc curvature.
To convert from effective distance to absolute screen distance, we used the distance to the pulsar of $372\pm 3{\rm\,pc}$ (Deller et al., 2019). The same measurements also gave an accurate proper motion, of $(-73.78,366.57){\rm\,mas/yr}$ in right ascension and declination, respectively, which, combined with the average Earth velocity during the observation (calculated using Astropy; Astropy Collaboration et al. 2022) allowed to infer the velocity of the screen from the effective velocity.
@@
Arecibo-VLA, $\Delta t_{V}~{}(\rm s)$1.5±0.8 -8.1±0.7 -18.3±2.3
  Arecibo-VLA, $\Delta t_{I}~{}(\rm s)$3.6±1.7 -9.9±0.5 -18.5±1.3
Arecibo-Jodrell Bank, $\Delta t_{V}~{}(\rm s)$3.7±1.0 -3.4±0.6 4.5±1.4
  Arecibo-Jodrell Bank, $\Delta t_{I}~{}(\rm s)$5.8±1.5 -4.6±0.6 4.1±1.1
Arecibo-Effelsberg, $\Delta t_{V}~{}(\rm s)$4.0±1.1 -2.7±0.6 6.9±2.2
  Arecibo-Effelsberg, $\Delta t_{I}~{}(\rm s)$ 2.5±2.9 -4.1±1.1 9.9±3.8
Arecibo-Westerbork, $\Delta t_{V}~{}(\rm s)$4.9±2.1 -3.7±1.0 6.2±3.2
  Arecibo-Westerbork, $\Delta t_{I}~{}(\rm s)$4.0±4.7 -8.9±2.8 8.7±2.3
@@
Line of Images, $\alpha_{s}$ (E of N) -100±030 -66±03 -37±03
Effective Velocity, $v_{{\rm eff},\parallel}~{}\rm(km/s)$ 1200±400 440±40 200±30
Effective Distance, $d_{\rm eff}~{}\rm(pc)$ 1100±900 350±80 240±80
Screen Velocity, $v_{\rm screen}~{}\rm(km/s)$200±300 -50±40 90±50
Screen Distance, $d_{\rm screen}~{}\rm(pc)$ 280±050 180±20 140±30
Figure 4: Baselines and scintillation screen orientations, in the $u-v$ plane. The blue, purple, grey and black points represent the Arecibo-VLA, Arecibo-Jodrell Bank, Arecibo-Effelsberg, and Arecibo-Westerbork baselines over the duration of the observation. The line from the origin to the mean location of each baseline represents the value we used in calculating the screen parameters. The lines surrounded by shaded regions represent the orientations of the lines of images and their uncertainties for each of the three screens, as inferred from the time delays. As the pulsar moves behind the screens in the direction indicated with the black arrow, the scintillation pattern for each screen moves across the sky, with a velocity vector opposite to that of the projection of the pulsar velocity on the screen.
@@
Perhaps surprising is that the nominal screen velocities we find are relatively large, $\gtrsim\!50{\rm\,km/s}$, well in excess of the expected velocities of $\sim\!7{\rm\,km/s}$ for gas co-rotating with the Galaxy using the model from Bovy (2017). This might be reflecting that Galactic rotation models are known to be poor approximations at high Galactic latitudes. However, all of the screen velocities have large error bars and are consistent with $0{\rm\,km/s}$ at the 1 to $2\sigma$ level. The large uncertainties are a result of large, positive covariance between the screen’s velocity and its distance — which in turn is because they both depend on the effective distance, which is relatively poorly constrained as our arc curvature measurements are not very precise. If the true screen velocities are closer to zero, then the distances would be somewhat smaller than those listed in Table 1 (given the near-unity covariance, by the same number of $\sigma$ by which the velocity deviates from 0, i.e., by $(\sigma_{d}/\sigma_{v})|v_{\rm screen}|$).

### v2

@@ Abstract
The scintillation of pulsars reveals small-scale structure of the interstellar medium. A powerful technique for characterizing the scintillating structures (screens) combines analysis of scintillation arcs and very long baseline interferometry (VLBI). We present the results of a VLBI analysis of the scintillation arcs of PSR B1133+16 from simultaneous observations with Arecibo, VLA, Jodrell Bank, Effelsberg, and Westerbork. Three arcs appear in the data set, all of which appear consistent with being the result of very anisotropic scattering screens. We are able to measure their orientations on the sky, down to uncertainties of $5\arcdeg$ for the two stronger screens, and measure distances, of $140\pm 30$, $180\pm 20$, and $250\pm 30{\rm\,pc}$, consistent with, but substantially more precise than what was inferred previously from annual modulation patterns in the scintillation. Comparing with the differential dust extinction with distance in this direction, the two nearer screens appear associated with the wall of the Local Bubble.
@@ 2.2 Visibility Secondary Cross-Spectra
$\frac{d\Phi_{S_{V}}}{df_{D}}=-\frac{2\pi|\mathbf{b}_{mn}|\cos(\alpha_{mn}-\alpha_{s})}{v_{\textrm{eff},\parallel}}=2\pi\Delta t_{V,mn},$ (8)
where $\alpha_{mn}$ and $\alpha_{s}$ represent the position angles of the baseline $\mathbf{b}_{mn}$ and the scattered images in the $(u,v)$ plane, respectively, and we implicitly defined $\Delta t_{V,mn}$, the time delay between when the scintillation pattern arrives at station $n$ compared to station $m$. One sees that the delay is maximized when the baseline aligns with the line of images, while it becomes zero if it is perpendicular to it: at that orientation, for a strictly linear set of images, the telescopes see identical interference patterns.
@@ 2.3 Intensity Secondary Cross-Spectra
$\frac{d\Phi_{S_{I}}}{df_{D}}=-\frac{2\pi|\mathbf{b}_{mn}|\cos(\alpha_{mn}-\alpha_{s})}{v_{\textrm{eff},\parallel}}=2\pi\Delta t_{I,mn}.$ (11)
@@ 2.4 Solving for Screen Parameters
A scintillation screen producing a (nearly) one-dimensional set of scattering images is described by three parameters, $d_{\rm eff}$, $v_{{\rm eff},\parallel}$, and $\alpha_{s}$. These three enter measurables via the curvature $\eta$ and the time delays $\Delta t_{V/I,mn}$, i.e., via Equations 6, 8, and 11 — all other parameters in those equations are known. Specifically, with at least two time delays (on non-parallel baselines), one can solve for $v_{{\rm eff},\parallel}$ and $\alpha_{s}$, and then combine $v_{{\rm eff},\parallel}$ with the curvature to find $d_{\rm eff}$.
@@ 3 Observations and Data Reduction
We analyze 2 hours of European VLBI Network (EVN) observations of PSR B1133+16 taken on 2019 January 27 (MJD 58510) from 6:29 to 8:29 UTC (PI: Kirsten, project code: GK049E). Data were taken with Arecibo, the Very Large Array (VLA), Jodrell Bank, Effelsberg, Westerbork and Sardinia, but we had to exclude the Sardinia data as these contained no pulsar signal likely due to issues with the telescope. For phase calibration of the VLA, the observations were interspersed every 26 minutes with a one-minute observation of the source J1143+221. A bandpass calibrator, J1125+2610, was observed by all telescopes using two five-minute scans separated by 80 minutes, but ended up unused in our analysis (see Section 3.1). The frequency range covered was from 316 MHz to 348 MHz, in two contiguous 16 MHz bands. All telescopes used circular polarizations except for VLA, which was linear. The dual-polarization baseband data were recorded in 2 bit real values, using VDIF format at all telescopes but Arecibo, where Mark 5B was used.
@@ 3.1 Reduction of Visibilities
Calibration of the data was done using CASA (McMullin et al., 2007). For our analysis here, which focuses on the properties of the screens, phase and polarization information is not relevant, so we averaged the eight on-pulse bins together and discarded the cross-polarizations. We flagged bad data and radio frequency interference, combining the flag table output from the EVN pipeline with manual inspection, and then used standard CASA routines for a-priori gain calibration, fringe fitting, bandpass calibration, delay calibration, and amplitude and phase corrections. We did not perform flux density scale calibration as the absolute flux is not relevant for scintillation measurements. As a significant portion of the bandpass calibrator observations were flagged or did not include all antennas, we found that the bandpass was flattened more successfully when using the pulsar for calibration directly instead of the calibrator source.
@@ 3.2 Reduction of Intensities
[…]
@@
Figure 2: Visibility secondary cross-spectra. Top: Phases and intensities for all baselines involving Arecibo (binned by a factor of 4 in $f_{\textrm{D}}$ and 8 in $\tau$). Parabolae for each arc for the best-fit arc curvature are overdrawn, with colours following the panels below. Bottom: Phase along the arc, determined by averaging the complex values along delay within a parabolic region around each arc of width $0.58{\rm\,mHz}$. Grey points indicate all measured values, while black ones indicate the ones used for fitting gradients. The best fit is shown with coloured lines, with the uncertainties represented by the surrounding shaded regions.
@@ 4 Scintillation Properties
In Figures 2 and 3, we show the visibility and intensity secondary cross-spectra, with color representing phase. One sees that on the Arecibo-VLA baseline, the inner and middle arcs have similar phase behavior, while on the Arecibo-Jodrell Bank baseline, the phases change in opposite directions along the parabola, showing immediately that these two screens have different orientations. We measure the time delays corresponding to the phase gradients in Section 4.2.
@@ 4.1 Arc Curvatures
While arcs can often be easily discerned, it is non-trivial to determine their curvatures. This is because any given image in an arc produces not just an interference signal with the line of sight, but also with every other image. Hence, except in the rare cases that images are very weak or distributed exactly symmetrically around the line of sight, the intensity-weighted center of an arc at a given delay will be offset from where it would be if only interference with the line of sight mattered.
In our case, as can be seen in Figure 1), power asymmetry is indeed present in the arms of the arcs, with the outer and inner arcs appearing brighter at negative Doppler values, and the middle arc appearing brighter at positive Doppler values, and this asymmetry appears to be reflected in the arclets (which are most noticeable in the inner arc). Furthermore, the analysis is hindered by the fact that the arcs overlap. While we describe below a quantative method that we believe gives good estimates of the curvatures, we note that the precise values are unpleasantly dependent on parameter choices. We attempt to account for these difficulties by relying on visual inspection both to confirm the estimates and to ensure the uncertainties are suitably conservative.
Figure 4: The mean power along parabolae of different curvatures of the Arecibo secondary spectrum (grey lines) using a range of bounds in doppler and delay. The vertical lines and shaded regions indicate the best fits and errors for the curvatures for the outer (green), middle (orange) and outer (blue) arcs.
For our estimates of the arc curvatures, we averaged the power along parabolae of different curvatures and identified peaks in the mean power of the Arecibo secondary spectrum. The widths of parabolae were set as 0.28 mHz (two pixels), and we intergrate only in the region between two ellipses, given by $(\tau/\tau_{\rm min})^{2}+(f_{\rm{D}}/f_{\rm D,min})^{2}=1$ and $(\tau/\tau_{\rm max})^{2}+(f_{\rm{D}}/f_{\rm D,max})^{2}=1$, to exclude the bright center where the arcs overlap as well as the noisy outskirts that carry little signal. The values of $\tau_{\rm min}$ were varied from 18 to 22 $\rm\mu s$, $\tau_{\rm max}$ from 30 to 50 $\rm\mu s$, $f_{\rm D,min}$ from 15 to 25 mHz, and $f_{\rm D,max}$ from 25 to 35 mHz.
Inspecting the result, shown in Fig. 4, one sees that the distribution of power shows a well-defined peak for the middle arc, and clear but less well-defined bumps with plateaus for the inner and outer ones. The effect of the asymmetry of brightness of arclets can be seen as the brighter bump at higher curvatures for the inner arc. Overlaying parabolae with curvatures centred on these peaks on the secondary spectrum, they seem good estimates of the arc curvatures (e.g., for the inner arc, they intersect the apexes of the arclets), so we take these as our best estimates. For the uncertainties, we fit three Gaussians to the power distribution for each set of ($\tau_{\rm min}$ $\tau_{\rm max}$$f_{\rm D,min}$$f_{\rm D,max}$) and use the range of values found for the means of the Gaussian as the error in the arc curvatures. We list the results in Table 1.
@@ 4.2 Phase Gradients
To estimate the time delays along each baseline, we measured the phase gradients as a function of Doppler in the visibility and intensity secondary cross-spectra (see Eqs. 8 and 11). First, we averaged the complex secondary cross-spectra over 4 and 8 bins in Doppler and delay, respectively, i.e., to a resolution of $0.56{\rm\,mHz}$ and $0.25{\rm\,\mu s}$. Then, for each arc, we averaged the complex values along the delay axis within a parabolic mask – centered around the best-fit parabola for the given arc and with a small width, $0.58{\rm\,mHz}$ – and determined the phases from those. For the uncertainties of the complex averages, we used the error in the mean, with the noise in the individual points estimated from the spectra far away from the arcs. We show the delay-averaged phases for each arc and baseline together with the binned secondary cross-spectra in Figures 2 and 3.
@@
We tested the influence of our assumptions on our fitted values by repeating the fits using a variety of parabolic masks and fitting regions. Specifically, we varied the arc curvature for the masks within the arc curvature fitting error, the widths from 0.56 to 1.12 mHz, and the minimum and maximum Doppler ranges by $\pm 2{\rm\,mHz}$ from those listed above. The fits depended most strongly on the arc curvature of the parabolic mask and had very little dependence on the widths of the parabolic mask or on the Doppler ranges. Overall, though, the fits deviated from each other beyond the nominal fitting error, so we take the standard deviation among the trials as the estimate of our uncertainty. We list the results in Table 1.
@@ 4.3 Phase De-Rotation
As an alternative method to determine the phase gradients in the secondary cross-spectra, we applied phase gradients of opposite sign directly to the secondary cross-spectra, i.e. multiplied by $\exp(-\Phi/f_{\textrm{D}})$. If the applied phase gradient is close to the true phase gradient, the imaginary component of the secondary cross-spectrum vanishes or, equivalently, the real component is maximized. We thus tried to find the best-fit phase gradient by first computing the sum of real values within the same masked regions as above for several test phase gradients, and then determining the location of the maximum by fitting a parabola to the results. This method gave consistent results to fitting the slope of the phases with Doppler.
@@ 5 Screen Properties
The screen parameters are found by using Equations 8 and 11 to infer from the time delays the orientation of the line of images and the effective velocity for each screen. For the inner and middle screen we fit all corresponding time delays simultaneously, derived from both intensity and visibility secondary cross-spectra, for the baselines between Arecibo and the four other antennas. For the outer arc, we included only the visibility secondary cross-spectra, as the fitting time delays for the intensity secondary cross-spectra were unreliable for the Arecibo-Effelsberg and Arecibo-Westerbork baselines. This is likely due to large amounts of radio frequency interference in the Effelsberg and Westerbork observations (which does not much affect the visibilities). Next, we use Equation 6 to calculate the effective distance from the effective velocity and the measured arc curvature for each screen.
To convert from effective distance to absolute screen distance, we used the distance to the pulsar of $372\pm 3{\rm\,pc}$ (Deller et al., 2019). The same measurements also gave an accurate proper motion, of $(-73.785^{+0.031}_{-0.010},366.569^{+0.072}_{-0.055}){\rm\,mas/yr}$ in right ascension and declination, respectively, which, combined with the average Earth velocity during the observation (calculated using Astropy; Astropy Collaboration et al. 2022) allowed us to infer the velocity of the screen from the effective velocity.
@@
Arecibo-VLA, $\Delta t_{V}~(\rm s)$. -2.2±0.5 -8.3±0.6 -18.3±1.8
  Arecibo-VLA, $\Delta t_{I}~(\rm s)$. -4.2±1.3 -9.9±0.5 -18.5±1.1
Arecibo-Jodrell Bank, $\Delta t_{V}~(\rm s)$. -3.8±1.0 -3.6±0.5 4.5±1.1
  Arecibo-Jodrell Bank, $\Delta t_{I}~(\rm s)$. -6.2±1.3 -4.4±0.6 4.1±0.9
Arecibo-Effelsberg, $\Delta t_{V}~(\rm s)$. -5.7±1.3 -3.0±0.6 6.9±1.9
  Arecibo-Effelsberg, $\Delta t_{I}~(\rm s)$. 0.4±3.4 -2.5±1.1 9.9±2.3
Arecibo-Westerbork, $\Delta t_{V}~(\rm s)$. -7.9±1.9 -3.5±1.0 6.2±3.0
  Arecibo-Westerbork, $\Delta t_{I}~(\rm s)$. 3.5±4.3 -8.3±3.2 8.7±3.3
@@
Line of Images, $\alpha_{s}$ (E of N). 66±011 115±05 143±04
Effective Velocity, $v_{{\rm eff},\parallel}~\rm(km/s)$. 1010±170 430±50 200±30
Effective Distance, $d_{\rm eff}~\rm(pc)$. 800±300 350±90 240±70
Screen Velocity, $v_{\rm screen}~\rm(km/s)$. -230±90 50±40 -90±40
Screen Distance, $d_{\rm screen}~\rm(pc)$. 250±030 180±20 140±30
Figure 5: Baselines and scintillation screen orientations, in the uvu-v plane. The blue, purple, grey and black points represent the Arecibo-VLA, Arecibo-Jodrell Bank, Arecibo-Effelsberg, and Arecibo-Westerbork baselines over the duration of the observation. The line from the origin to the mean location of each baseline represents the value we used in calculating the screen parameters. The lines surrounded by shaded regions represent the orientations of the lines of images and their uncertainties for each of the three screens, as inferred from the time delays. As the pulsar moves behind the screens in the direction indicated with the black arrow, the scintillation pattern for each screen moves across the sky, with a velocity vector opposite to that of the projection of the pulsar velocity on the screen.
@@
Perhaps surprising is that the nominal screen velocities we find are relatively large, $\gtrsim\!50{\rm\,km/s}$, well in excess of the expected velocities of $\sim\!7{\rm\,km/s}$ for gas co-rotating with the Galaxy using the model from Bovy (2017). This might be reflecting that Galactic rotation models are known to be poor approximations at high Galactic latitudes. However, all of the screen velocities have large error bars and are consistent with $0{\rm\,km/s}$ at the 1.5 to $3\sigma$ level. The large uncertainties are a result of large, positive covariance between the screen’s velocity and its distance — which in turn is because they both depend on the effective distance, which is relatively poorly constrained as our arc curvature measurements are not very precise. If the true screen velocities are closer to zero, then the distances would be somewhat smaller than those listed in Table 1 (given the near-unity covariance, by the same number of $\sigma$ by which the velocity deviates from 0, i.e., by $(\sigma_{d}/\sigma_{v})|v_{\rm screen}|$).